OpenStudy (anonymous):

A car starts from rest at a stop sign. It accelerates at 4.0 m/s2 for 9 seconds, coasts for 2 s, and then slows down at a rate of 3.0 m/s2 for the next stop sign. How far apart are the stop signs?

OpenStudy (anonymous):

Acceleration phase: Vf = Vi + a*t Vf = 0 + (4.0)(9) = 36 m/s Vf^2 = Vi^2 + 2a(Xf - Xi) 36^2 = 0 + 2(4.0)(Xf - 0) 1296 = 8 Xf Xf = 162 m Coasting phase: v = 36 m/s t = 2 s Xf = v * t = (36)(2) = 72 m Deceleration phase: Vf^2 = Vi^2 + 2a(Xf - Xi) 0 = 36^2 + 2(-3.0)(Xf - 0) -1296 = -6 Xf Xf = 216 m Find the total distance: Xf acceleration = 162 m Xf coasting = 72 m Xf deceleration = 216 m Sum of these: Xtotal = 162 + 72 + 216 = 450 m

OpenStudy (anonymous):

thanks so much!